Suppose that X1,... , X are i.i.d. with an exponential distribution with mean equal to 2.0. Assume that n = 40. Let X, = 1-1Xapproximate distribution for 1/Xn, using the delta method Obtain an
Answer #1

Answer:- PI Given that > n = 40 consider In=1&hxa. Now we have to assume that. Xin mexn Gluen that Mian: 20 Consider the EquaPg Ex - 8 .. Var (x;J = E(2;2- E* () 7 = 8-(2) Tilhere Elad) -2, 78-4 L Hi-Ion 4 Hence . So Var (21) - H Ye; is 12,...n HeneNow we have to calculate approui ma te Pg@ distribution for I By the above Solution. here Eī <40 . [12] 40 il 00 I IPH __ 1 4pg Hence ? . Var[In] un * Varlton] 40 E vai(78) 40 lat i so 4062 4-02 402 4062 цохиб 4062 40 Where to ² - 4 = 40 = 10PgG aproximata. We have find the Solution distribution for z Where most Centras limit theorem Here we have useck to nie ding